OpenStudy (anonymous):

Use the Table to evaluate each expression. x | 1 2 3 4 5 6 f | 3 1 4 2 2 5 g | 6 3 2 1 2 3 a). f(g(1)) b). g(f(1)) c). f(f(1)) d). g(g(1)) e). (gof)(3) f). (fog)(6) help me understand the concept and how to do it, and I'll try on my own

OpenStudy (helder_edwin):

following your table u got \[ \large f(g(1))=f(6)=5 \] got it???

OpenStudy (anonymous):

nope.. I need like step by step instructions.. ?

OpenStudy (ghazi):

g(1) means value of g at x=1

OpenStudy (helder_edwin):

ok. but u do understand the table, don't u???

OpenStudy (anonymous):

I don't know how to use the table to find the answer. That's my issue.

OpenStudy (ghazi):

now g at x=1 is 6 ...therefore now F(6) is the value of f at x=6 which is 5

OpenStudy (helder_edwin):

the first line gives u the value of the independent variable: 1 through 6 the second line tells u that the function \(f\) transforms 1 into 3 hence f(1)=3 got it??

OpenStudy (anonymous):

okay, so for the second one.. x=1 f(3) g(2).. so 2 is the answer?

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

and the third is 3?

OpenStudy (helder_edwin):

\[ \large g(f(1))=g(3)=2 \]

OpenStudy (helder_edwin):

try writing it the way i did!!!

OpenStudy (helder_edwin):

no, the third is not 3

OpenStudy (anonymous):

oh four?

OpenStudy (anonymous):

f(f(1))=f(3)=4

OpenStudy (helder_edwin):

yes. u got it!!

OpenStudy (anonymous):

g(g(1))=g(6)=3?

OpenStudy (helder_edwin):

yes

OpenStudy (helder_edwin):

\[ \large (g\circ f)(3)=g(f(3))= \]

OpenStudy (anonymous):

how do I solve the last two cause they're differentt...

OpenStudy (anonymous):

is (gof)(3)=1? (fog)(6)=4?

OpenStudy (helder_edwin):

yes to both!! did u understand???

OpenStudy (anonymous):

yeah thank you :)

OpenStudy (helder_edwin):

u r welcome